Net sales = $42,000,000. Gross profit = $19,200,000. Other operating expenses = $3,200,000. Addition to retained earnings = $5,200,000. Dividends paid to preferred and common stockholders = $2,000,000. Depreciation expense = $2,930,000.

The firm’s tax rate is 36 percent. interest expense for Moore’s HoneyBee Corp.
Answer #1

Net Sales COGS Gross Profit Other Operating Expenses EBITDA Depreciation & Amortizatio EBIT Interest EBT Tax %-36% PAT 42000000 22800000 19200000 3200000 16000000 2930000 13070000 1820000 11250000 4050000 7200000 Amortization assumed =0 as not mentioned

PAT = Retained Earnings + Dividends Paid = $5,200,000 + $2,000,000 = $7,200,000
Since Tax = 36% therefore PAT = 64% of EBT/PBT = Earnings/Profit Before Tax
Thus we calculated EBT = PAT / 0.64 = $7,200,000 / 0.64 = $11,250,000
We get EBIT as $13,070,000 after deduction of Depreciation.

Thus EBIT - Interest Expense = PBT/EBT.

Thus we calculated Interest Expense = EBIT - EBT = $13,070,000 - $11,250,000 = $1,820,000